OpenStudy (anonymous):

A charity is selling raffle tickets at $5 apiece. Out of the 100 tickets sold, ten third-place winners will be drawn, each winning $20. Five second-place winners will be drawn, each winning $50. Lastly, one grand prize of $100 will be awarded. Is a raffle with these rules a good idea for the charity? Why or why not?

jimthompson5910 (jim_thompson5910):

how far did you get with this @kota112 ?

OpenStudy (anonymous):

i know its yes im

jimthompson5910 (jim_thompson5910):

why do you think "yes" ?

OpenStudy (anonymous):

because i know its worth 50 cents

OpenStudy (anonymous):

im trying to find out if its more or less than face value

jimthompson5910 (jim_thompson5910):

50 cents? where are you getting that?

OpenStudy (dumbcow):

i know im buying these raffle tickets

OpenStudy (dumbcow):

Charity gets 5*100 = 500 Charity pays out winnings ---> 10*20 + 5*50 + 1*100 = 550 Charity loses 50 dollars

OpenStudy (anonymous):

so its c?

jimthompson5910 (jim_thompson5910):

you didn't list any answer choices

OpenStudy (anonymous):

oh sorry i was looking at the wrong question

OpenStudy (dumbcow):

lol ??

OpenStudy (anonymous):

heres the options haha just made a new account still getting used to it

OpenStudy (anonymous):

No, because one ticket’s expected value is worth $0.50 more than the face value. B. Yes, because one ticket’s expected value is worth $0.50 more than the face value. C. Yes, because one ticket’s expected value is worth $0.50 less than the face value. D. No, because one ticket’s expected value is worth $0.50 less than the face value

jimthompson5910 (jim_thompson5910):

dumbcow has the right idea overall the charity will lose $50 if anything, they will probably pay back the $500 and not be able to pay that additional $50

jimthompson5910 (jim_thompson5910):

@kota112 what should the price of the ticket be so that it's a fair game? By "fair game" I mean that the charity gains no money nor loses money.

OpenStudy (anonymous):

they will lose money

jimthompson5910 (jim_thompson5910):

Let p = price of each ticket As dumbcow pointed out, the charity will pay out $550 based on the rules above They will get 100p dollars coming in If it's a fair game, then the difference of the revenue and expenses will be zero 550 - 100p = 0 solve for p to get p = ??

jimthompson5910 (jim_thompson5910):

oops I had that backwards, it should be 100p - 550 = 0 but in this case, the value of p remains the same

jimthompson5910 (jim_thompson5910):

anyways, what do you get when you solve for p?

OpenStudy (anonymous):

.50 less than face value so it would be d because its not a good idea for charity

jimthompson5910 (jim_thompson5910):

well if you are in the person's shoes who wins the money then the expected value of their winnings is $5.50 which is 50 cents more than the ticket price ($5) so on average, the charity is losing 50 cents per ticket multiply that by 100 and you get 100*(50 cents) = 50 dollars total loss

OpenStudy (anonymous):

okay thanks a lot:)

jimthompson5910 (jim_thompson5910):

np